OpenStudy (anonymous):

how would i find the exact value of the square root of 2 over the square root of 8?

OpenStudy (zzr0ck3r):

sqrt(2)/sqrt(8) = sqrt(2/8) = sqrt(1/4) = 1/sqrt(4) = 1/2